OpenStudy (anonymous):

Indicate which property is illustrated in Step 2. Step 1 7 - 8x - 7 + 9x = (7 - 8x) + (-7 + 9x) Step 2 = (7 - 8x) + (9x - 7) Step 3 = 7 + (-8x + 9x) - 7 Step 4 = 7 + (-8 + 9)x - 7 Step 5 = 7 + x - 7 Step 6 = x + 7 - 7 Step 7 = x + 0 Step 8 = x identity distributive associative commutative

Directrix (directrix):

What did you put?

OpenStudy (anonymous):

associative

Directrix (directrix):

(7 - 8x) + (-7 + 9x) to (7 - 8x) + (9x - 7)

Directrix (directrix):

(7 - 8x) + (-7 + 9x) to (7 - 8x) + (9x - 7)

Directrix (directrix):

This (-7 + 9x) to (9x - 7) The order has changed. Which property? NOT associative.

OpenStudy (anonymous):

starts with a d

Directrix (directrix):

starts with a c

Directrix (directrix):

@Starr_Dynasty

OpenStudy (anonymous):

Indicate which property is illustrated in Step 1. Step 1 7x + 11 + 2x + 4 = (7x + 11) + (2x + 4) Step 2 = 7x + (11 + 2x) + 4 Step 3 = 7x + (2x + 11) + 4 Step 4 = (7x + 2x) + (11 + 4) Step 5 = (7 + 2)x + (11 + 4) Step 6 = 9x + 15 distributive associative arithmetic fact commutative

Directrix (directrix):

It is the grouping property again.

OpenStudy (anonymous):

associative

Directrix (directrix):

Correct

OpenStudy (anonymous):

Indicate which property is illustrated in Step 3. Step 1 3x + 11 + 9x + 9 = (3x + 11) + (9x + 9) Step 2 = 3x + (11 + 9x) + 9 Step 3 = 3x + (9x + 11) + 9 Step 4 = (3x + 9x) + (11 + 9) Step 5 = (3 + 9)x + (11 + 9) Step 6 = 12x + 20 distributive associative commutative arithmetic fact

Directrix (directrix):

@Starr_Dynasty What did you get?

OpenStudy (anonymous):

2nd

Directrix (directrix):

(11 + 9x) to (9x + 11) NOT Associative Ordering Property

Directrix (directrix):

See chart to review.

OpenStudy (anonymous):

Indicate which property is illustrated in Step 4. Step 1 9 - 8x - 9 + 9x = (9 - 8x) + (-9 + 9x) Step 2 = (9 - 8x) + (9x - 9) Step 3 = 9 + (-8x + 9x) - 9 Step 4 = 9 + (-8 + 9)x - 9 Step 5 = 9 + x - 9 Step 6 = x + 9 - 9 Step 7 = x + 0 Step 8 = x distributive associative commutative identity

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

After this problem, start a new thread, please.

Directrix (directrix):

9 + (-8x + 9x) - 9 to 9 + (-8 + 9)x - 9

Directrix (directrix):

This is what has changed: (-8x + 9x) to (-8 + 9)x That's our friend with a "D."

Directrix (directrix):

@Starr_Dynasty ^^^^

Directrix (directrix):

END OF THIS THREAD --- PLEASE START NEW ONE